A youth organization collected nickels and dimes for a charity drive. by the end of the 1st day s drive, the youth had collected 56.35. If there were three times as many dimes as nickels, how many of each type of coin was collected?
Please Help

Answer by JimboP1977(308)

Dime = 0.1 = D
Nickel = 0.05 = N
D:N = 3:1 = 0.30 : 0.05
So we have a certain unknown number multiplied by 3 lots of Dimes plus 1 lot of Nickels equals 56.35 or using algebra:
x * (0.30 + 0.05) = 56.35
We need to know what x is so we rearrange by dividing both sides by 0.35 which gives x as 161.
This tells us that 161 coins were nickels and three times this amount were dimes. So 483 were dimes.
To double check:
161 * 0.05 = $8.05
483 * 0.10 = $48.30
Total = $56.35
Does this help?